# Scaling limit of the staggered six-vertex model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant boundary conditions

### v2

@@ 4.2 ${\cal W}_{\infty}$ algebra
$\bigoplus_{{\tt d}=0}^{\infty}{\cal H}_{p,s}^{({\tt d})}\,,\qquad\qquad{\cal H}_{p,s}^{({\tt d})}=\begin{cases}{\cal H}_{p,s}^{({\rm cont},{\tt d})}&\ \ \ {\rm for}\qquad s\in\mathbb{R}\\[5.69046pt] {\cal H}_{p,s}^{({\rm disc},{\tt d})}&\ \ \ {\rm for}\qquad p+\tfrac{1}{2}\pm{\rm i}s\in\mathbb{Z}\end{cases}$ (4.24)
@@
$\widetilde{W}_{j}(m)\,|\bm{\omega}\rangle=0\qquad\qquad(\forall m>0)\,,\qquad\qquad\widetilde{W}_{j}(0)\,|\bm{\omega}\rangle=\omega_{j}\,|\bm{\omega}\rangle$ (4.31)
with $j=2,3$. The highest weight is given by $\bm{\omega}=(\omega_{2},\omega_{3})$, where the component $\omega_{2}$ is equal to the conformal dimension of the highest state, while $\omega_{3}$ is the eigenvalue of $\widetilde{W}_{3}(0)$, which commutes with $\widetilde{W}_{2}(0)$. The Verma module is spanned by the states that are obtained by acting with the ‘creation modes’ of the spin 2 and spin 3 currents on the highest state:
$\widetilde{W}_{2}(-\ell_{1})\ldots\widetilde{W}_{2}(-\ell_{m})\,\widetilde{W}_{3}(-\ell_{1}^{\prime})\ldots\widetilde{W}_{3}(-\ell_{m^{\prime}}^{\prime})|\bm{\omega}\rangle$ (4.32)
@@
${\tt d}=\sum_{j=1}^{m}\ell_{j}+\sum_{j=1}^{m^{\prime}}\ell_{j}^{\prime}$ (4.33)
and the dimensions of its level subspace with fixed ${\tt d}$ is the number of bi-partitions of ${\tt d}$, i.e., ${\rm par}_{2}({\tt d})$ (3.28). In what follows we will parameterize the highest weight for the Verma module ${\cal V}_{\rho,\nu}$ as
@@
This is motivated by the free field realization (4.28). Supposing that the highest state is an eigenvector of the operators $\int\mbox{d}u\,\partial\vartheta(u)$ and $\int\mbox{d}u\,\partial\varphi(u)$ with eigenvalues $\frac{\nu}{\sqrt{n}}$ and $\frac{\rho}{\sqrt{n+2}}$, respectively, formula (4.34) follows from (4.28). The highest weight is an even function of $\rho$. As a result the spaces ${\cal V}_{\rho,\nu}$ and ${\cal V}_{-\rho,\nu}$ should be identified. In the parameterization (4.34), the conformal dimensions of a state in the Verma module at level ${\tt d}$ is such that
$\widetilde{W}_{2}(0)-\frac{c}{24}=\frac{\rho^{2}}{n+2}+\frac{\nu^{2}}{n}-\frac{1}{12}+{\tt d}\,,$ (4.35)
@@
For generic complex values of $\rho$ and $\nu$ the Verma module ${\cal V}_{\rho,\nu}$ is an irreducible representation of the ${\cal W}_{\infty}$ algebra. However, with $\rho$, $\nu$ obeying certain constraints, the Verma module contains null vectors – highest states occuring at non-zero levels. Then the highest weight irrep ${\cal W}_{\rho,\nu}$ can be obtained from ${\cal V}_{\rho,\nu}$ by factoring out all of the invariant subspace(s) generated by the null vector(s). In view of applications to the scaling limit of the lattice model of particular interest is when $\rho=\pm\frac{1}{2}\big{(}r-m(n+2)\big{)}$ with $r,m=1,2,\ldots$. In this case a null vector occurs at level ${\tt d}=mr$ and the Verma module splits into the direct sum of two representations, which are irreducible for generic $n$ and $\nu$:
@@
The space ${\cal W}_{\rho^{\prime},\nu}$ is isomorphic to the Verma module and the dimensions of its level subspaces is ${\rm par}_{2}({\tt d})$, while for ${\cal W}_{\rho,\nu}$, the level subspaces are ${\rm par}_{2}({\tt d})-{\rm par}_{2}({\tt d}-mr)$ dimensional. Consider again the components ${\cal H}^{({\rm cont},{\tt d})}_{p,s}$, which appear in the decomposition of the continuous spectrum of the space of states occuring in the scaling limit of the spin chain. Taking into account formulae (4.12) and (3.30) it is clear that
${\cal W}_{p,s}\cong\bigoplus_{{\tt d}=0}^{\infty}{\cal H}^{({\rm cont},{\tt d})}_{p,s}\qquad\qquad\qquad\qquad\big{(}2p=2{\cal S}+1-(n+2)\,,\ s\ {\rm-\,real}\big{)}\,.$ (4.37)
To describe the discrete spectrum in terms of irreps of the ${\cal W}_{\infty}$ algebra, it is necessary to analyze the case when $\nu$ is such that $\rho+\frac{1}{2}\pm{\rm i}\nu$ is an integer for some choice of the sign $\pm$. As explained in, e.g., section 16.2 of the work [15] the Verma module with $\rho+\frac{1}{2}+{\rm i}\nu=-a_{+}=0,\pm 1,\pm 2,\ldots$ contains a null vector at level $|a_{+}+\frac{1}{2}|+\frac{1}{2}$, while for $-\rho+\frac{1}{2}+{\rm i}\nu=-a_{-}=0,\pm 1,\pm 2,\ldots$ there is a null vector at level $|a_{-}+\frac{1}{2}|+\frac{1}{2}$. Assuming $\rho$ is generic for now the character of the irreducible representation,
${\rm ch}_{\rho,\nu}({\tt q})\equiv{\rm Tr}_{{\cal W}_{\rho,\nu}}\Big{[}\,{\tt q}^{\widetilde{W}_{2}(0)-\frac{c}{24}}\,\Big{]}\,,$ (4.38)
@@
[…]
If, in addition to $\nu$ being constrained as above, $\rho\to\frac{1}{2}\,(2{\cal S}-n-1)$ then the irrep ${\cal W}_{\rho,\nu}$ further breaks up into two irreducible representations. One of them is generated by the null vector which appears at level $12{\cal S}+$ 1 and has highest weights $(\rho^{\prime},\nu)$ with $\rho^{\prime}=\frac{1}{2}\,(2{\cal S}+n+3)$. Its character is given by (4.39) with $\rho$ replaced by $\rho^{\prime}$ and $a\mapsto a^{\prime}=-2({\cal S}+1)-a$. Taking the difference ${\rm ch}_{\rho,\nu}({\tt q})-{\rm ch}_{\rho^{\prime},\nu}({\tt q})$ with $\rho\to\frac{1}{2}\,(2{\cal S}-n-1)$ and $\rho^{\prime}\to\frac{1}{2}\,(2{\cal S}+n+3)$ yields for the character of the irreducible representation ${\cal W}_{\rho,\nu}$ with
@@
${\rm ch}_{\rho,\nu}={\tt q}^{-\frac{1}{12}+\frac{\nu^{2}}{n}+\frac{\rho^{2}}{n+2}}\ \bigg{(}\prod_{j=1}^{\infty}\frac{1}{(1-{\tt q}^{j})^{2}}\bigg{)}\sum_{m=0}^{\infty}(-1)^{m}\,q^{\frac{m^{2}}{2}}\,\Big{(}q^{m|a+\frac{1}{2}|}-q^{2{\cal S}+1+m|2{\cal S}+a+\frac{3}{2}|}\Big{)}\ .$ (4.41)
For the case $a\geq 0$ the above expression, apart from the overall factor ${\tt q}^{-\frac{1}{12}+\frac{\nu^{2}}{n}+\frac{\rho^{2}}{n+2}}$, coincides with the generating function (4.7) for the integers $N_{a}^{+}({\tt d}\,|\,{\cal S})$. This way, one concludes
${\cal W}_{p,s}\cong\bigoplus_{{\tt d}=0}^{\infty}{\cal H}^{({\rm disc},{\tt d})}_{p,s}\qquad\qquad\qquad\big{(}p={\cal S}+\tfrac{1}{2}-\tfrac{1}{2}(n+2)\,,\ \tfrac{n}{2}-{\cal S}\pm{\rm i}s=a\in\mathbb{Z}_{\geq 0}\big{)}\,.$ (4.42)
The remaining case to be considered is when $-\mathcal{S}\leq a<0$. The lower bound comes from the condition $s\in(-\frac{n}{2},\frac{n}{2}]$ which implies that ±$\pm{\rm i}s=\tfrac{n}{2}-{\mathcal{S}}-a\leq\frac{n}{2}$. From the definition of the integers $N^{+}_{a}({\tt d}\,|\,{\cal S})$ (4.7), which give the dimensions of the level subspaces ${\cal H}_{p,s}^{({\rm disc},{\tt d})}\subset{\cal H}^{({\rm disc})}$, one finds
$\dim\big{(}{\cal H}_{p,s}^{({\tt d})}\big{)}=0\qquad{\rm for}\qquad{\tt d}=0,1,\ldots,|a|-1\qquad\qquad\big{(}-p-\tfrac{1}{2}\pm{\rm i}s=a\in\mathbb{Z}\,,\ -{\cal S}\leq a<0\big{)}$ (4.43)
@@
${\cal W}_{\rho,\nu}=\bigoplus_{{\tt d}=0}^{\infty}{\cal H}_{p,s}^{({\tt d})}\,,$ (4.45)
@@
$\rho=\mathcal{S}+\frac{1}{2}\,,\qquad\qquad\nu=\begin{cases}s-\frac{{\rm i}n}{2}&{\rm for}\ \ \ (-{\rm i}s)>0\\[5.69046pt] s+\frac{{\rm i}n}{2}&{\rm for}\ \ \ (-{\rm i}s)<0\end{cases}\qquad\qquad\big{(}\tfrac{n}{2}-{\cal S}\pm{\rm i}s=a\in\mathbb{Z}_{<0}\,,\ -{\cal S}\leq a<0\big{)}\,.$
@@
${\rm ch}_{\rho,\nu}({\tt q})={\tt q}^{-\frac{1}{12}+\frac{\nu^{2}}{n}+\frac{\rho^{2}}{n+2}}\ \bigg{(}\prod_{m=1}^{\infty}\frac{1}{(1-{\tt q}^{m})^{2}}\bigg{)}\ \sum_{m=0}^{\infty}(-1)^{m}\,{\tt q}^{\frac{m^{2}}{2}}\big{(}\,{\tt q}^{m|\,|\rho|-|\nu|\,|}-{\tt q}^{(m+1)(|\rho|+|\nu|+1)-\frac{1}{2}}\,\big{)}\,.$ (4.46)
One can check that the dimensions of the level subspaces, obtained by expanding ${\rm ch}_{\rho,\nu}({\tt q})$ in a series in ${\tt q}$, coincides with the integers $N^{+}_{a}({\tt d}|\,{\cal S})$ with $-{\cal S}\leq a<0$ and ${\tt d}=|a|,\,|a|+1,\,|a|+2,\ldots$
Finally, we mention that the states $|\psi_{p,s}^{(S^{z})}(\bm{w})\rangle\in{\cal W}_{\rho,\nu}$ appearing in the scaling limit of the Bethe states, see eq.(4.9), have an important interpretation. They are the simultaneous eigenstates of the so-called quantum AKNS integrable structure [50, 51]. The function $\mbox{e}^{\frac{{\rm i}}{2}\delta}$ (3.41) entering into the quantization condition coincides with the eigenvalue of a certain so-called reflection operator [52] computed on $|\psi_{p,s}^{(S^{z})}(\bm{w})\rangle$, see ref.[45] for details.

### v3

@@ 4.2 ${\cal W}_{\infty}$ algebra
$\bigoplus_{\texttt{d}=0}^{\infty}{\cal H}_{p,s}^{(\texttt{d})}\,,\qquad\qquad{\cal H}_{p,s}^{(\texttt{d})}=\begin{cases}{\cal H}_{p,s}^{({\rm cont},\texttt{d})}&\ \ \ {\rm for}\qquad s\in\mathbb{R}\\[5.69046pt] {\cal H}_{p,s}^{({\rm disc},\texttt{d})}&\ \ \ {\rm for}\qquad p+\tfrac{1}{2}\pm{\rm i}s\in\mathbb{Z}\end{cases}$ (4.24)
@@
$\widetilde{W}_{j}(m)\,|\boldsymbol{\omega}\rangle=0\qquad\qquad(\forall m>0)\,,\qquad\qquad\widetilde{W}_{j}(0)\,|\boldsymbol{\omega}\rangle=\omega_{j}\,|\boldsymbol{\omega}\rangle$ (4.31)
with $j=2,3$. The highest weight is given by $\boldsymbol{\omega}=(\omega_{2},\omega_{3})$, where the component $\omega_{2}$ is equal to the conformal dimension of the highest state, while $\omega_{3}$ is the eigenvalue of $\widetilde{W}_{3}(0)$, which commutes with $\widetilde{W}_{2}(0)$. The Verma module is spanned by the states that are obtained by acting with the ‘creation modes’ of the spin 2 and spin 3 currents on the highest state:
$\widetilde{W}_{2}(-\ell_{1})\ldots\widetilde{W}_{2}(-\ell_{m})\,\widetilde{W}_{3}(-\ell_{1}^{\prime})\ldots\widetilde{W}_{3}(-\ell_{m^{\prime}}^{\prime})|\boldsymbol{\omega}\rangle$ (4.32)
@@
$\texttt{d}=\sum_{j=1}^{m}\ell_{j}+\sum_{j=1}^{m^{\prime}}\ell_{j}^{\prime}$ (4.33)
and the dimensions of its level subspace with fixed d is the number of bi-partitions of d, i.e., ${\rm par}_{2}(\texttt{d})$ (3.29). In what follows we will parameterize the highest weight for the Verma module ${\cal V}_{\rho,\nu}$ as
@@
This is motivated by the free field realization (4.28). Supposing that the highest state is an eigenvector of the operators $\int\mbox{d}u\,\partial\vartheta(u)$ and $\int\mbox{d}u\,\partial\varphi(u)$ with eigenvalues $\frac{\nu}{\sqrt{n}}$ and $\frac{\rho}{\sqrt{n+2}}$, respectively, formula (4.34) follows from (4.28). The highest weight is an even function of $\rho$. As a result the spaces ${\cal V}_{\rho,\nu}$ and ${\cal V}_{-\rho,\nu}$ should be identified. In the parameterization (4.34), the conformal dimensions of a state in the Verma module at level d is such that
$\widetilde{W}_{2}(0)-\frac{c}{24}=\frac{\rho^{2}}{n+2}+\frac{\nu^{2}}{n}-\frac{1}{12}+\texttt{d}\,,$ (4.35)
@@
For generic complex values of $\rho$ and $\nu$ the Verma module ${\cal V}_{\rho,\nu}$ is an irreducible representation of the ${\cal W}_{\infty}$ algebra. However, with $\rho$, $\nu$ obeying certain constraints, the Verma module contains null vectors – highest states occurring at non-zero levels. Then the highest weight irrep ${\cal W}_{\rho,\nu}$ can be obtained from ${\cal V}_{\rho,\nu}$ by factoring out all of the invariant subspace(s) generated by the null vector(s). In view of applications to the scaling limit of the lattice model of particular interest is when $\rho=\pm\frac{1}{2}\big{(}r-m(n+2)\big{)}$ with $r,m=1,2,\ldots$. In this case a null vector occurs at level $\texttt{d}=mr$ and the Verma module splits into the direct sum of two representations, which are irreducible for generic $n$ and $\nu$:
@@
The space ${\cal W}_{\rho^{\prime},\nu}$ is isomorphic to the Verma module and the dimensions of its level subspaces is ${\rm par}_{2}(\texttt{d})$, while for ${\cal W}_{\rho,\nu}$, the level subspaces are ${\rm par}_{2}(\texttt{d})-{\rm par}_{2}(\texttt{d}-mr)$ dimensional. Consider again the components ${\cal H}^{({\rm cont},\texttt{d})}_{p,s}$, which appear in the decomposition of the continuous spectrum of the space of states occurring in the scaling limit of the spin chain. Taking into account formulae (4.12) and (3.31) it is clear that
${\cal W}_{p,s}\cong\bigoplus_{\texttt{d}=0}^{\infty}{\cal H}^{({\rm cont},\texttt{d})}_{p,s}\qquad\qquad\qquad\big{(}2p=2{\cal S}+1-(n+2)\,,\ s\ {\rm-\,real}\big{)}\,.$ (4.37)
To describe the discrete spectrum in terms of irreps of the ${\cal W}_{\infty}$ algebra, it is necessary to analyze the case when $\nu$ is such that $\rho+\frac{1}{2}\pm{\rm i}\nu$ is an integer for some choice of the sign $\pm$. As explained in, e.g., section 16.2 of reference [15] the Verma module with $\rho+\frac{1}{2}+{\rm i}\nu=-a_{+}=0,\pm 1,\pm 2,\ldots$ contains a null vector at level $|a_{+}+\frac{1}{2}|+\frac{1}{2}$, while for $-\rho+\frac{1}{2}+{\rm i}\nu=-a_{-}=0,\pm 1,\pm 2,\ldots$ there is a null vector at level $|a_{-}+\frac{1}{2}|+\frac{1}{2}$. Assuming $\rho$ is generic for now the character of the irreducible representation,
${\rm ch}_{\rho,\nu}(\texttt{q})\equiv{\rm Tr}_{{\cal W}_{\rho,\nu}}\Big{[}\,\texttt{q}^{\widetilde{W}_{2}(0)-\frac{c}{24}}\,\Big{]}\,,$ (4.38)
@@
${\rm ch}_{\rho,\nu}(\texttt{q})=\texttt{q}^{-\frac{1}{12}+\frac{\nu^{2}}{n}+\frac{\rho^{2}}{n+2}}\ \bigg{(}\prod_{m=1}^{\infty}\frac{1}{(1-\texttt{q}^{m})^{2}}\bigg{)}\sum_{m=0}^{\infty}(-1)^{m}\ \texttt{q}^{m|a+\frac{1}{2}|+\frac{m^{2}}{2}}\qquad\quad\begin{array}[]{l}\rho+\tfrac{1}{2}\pm{\rm i}\nu=-a\in\mathbb{Z}\\[5.69046pt] n,\,\rho\ \ {\rm generic}\end{array}\,.$ (4.39)
If, in addition to $\nu$ being constrained as above, $\rho\to\frac{1}{2}\,(2{\cal S}-n-1)$ then the irrep ${\cal W}_{\rho,\nu}$ further breaks up into two irreducible representations. One of them is generated by the null vector which appears at level $12{\cal S}+$ 1 and has highest weights $(\rho^{\prime},\nu)$ with $\rho^{\prime}=\frac{1}{2}\,(2{\cal S}+n+3)$. Its character is given by (4.39) with $\rho$ replaced by $\rho^{\prime}$ and $a\mapsto a^{\prime}=-2({\cal S}+1)-a$. Taking the difference ${\rm ch}_{\rho,\nu}(\texttt{q})-{\rm ch}_{\rho^{\prime},\nu}(\texttt{q})$ with $\rho\to\frac{1}{2}\,(2{\cal S}-n-1)$ and $\rho^{\prime}\to\frac{1}{2}\,(2{\cal S}+n+3)$ yields for the character of the irreducible representation ${\cal W}_{\rho,\nu}$ with
@@
${\rm ch}_{\rho,\nu}=\texttt{q}^{-\frac{1}{12}+\frac{\nu^{2}}{n}+\frac{\rho^{2}}{n+2}}\ \bigg{(}\prod_{j=1}^{\infty}\frac{1}{(1-\texttt{q}^{j})^{2}}\bigg{)}\sum_{m=0}^{\infty}(-1)^{m}\,q^{\frac{m^{2}}{2}}\,\Big{(}q^{m|a+\frac{1}{2}|}-q^{2{\cal S}+1+m|2{\cal S}+a+\frac{3}{2}|}\Big{)}\ .$ (4.41)
For the case $a\geq 0$ the above expression, apart from the overall factor $\texttt{q}^{-\frac{1}{12}+\frac{\nu^{2}}{n}+\frac{\rho^{2}}{n+2}}$, coincides with the generating function (4.7) for the integers $N_{a}^{+}(\texttt{d}\,|\,{\cal S})$. This way, one concludes
${\cal W}_{p,s}\cong\bigoplus_{\texttt{d}=0}^{\infty}{\cal H}^{({\rm disc},\texttt{d})}_{p,s}\qquad\qquad\big{(}p={\cal S}+\tfrac{1}{2}-\tfrac{1}{2}(n+2)\,,\ \tfrac{n}{2}-{\cal S}\pm{\rm i}s=a\in\mathbb{Z}_{\geq 0}\big{)}\,.$ (4.42)
The remaining case to be considered is when $-\mathcal{S}\leq a<0$. The lower bound comes from the condition $s\in(-\frac{n}{2},\frac{n}{2}]$ which implies that ±$\pm{\rm i}s=\tfrac{n}{2}-{\mathcal{S}}-a\leq\frac{n}{2}$. From the definition of the integers $N^{+}_{a}(\texttt{d}\,|\,{\cal S})$ (4.7), which give the dimensions of the level subspaces ${\cal H}_{p,s}^{({\rm disc},\texttt{d})}\subset{\cal H}^{({\rm disc})}$, one finds
$\dim\big{(}{\cal H}_{p,s}^{(\texttt{d})}\big{)}=0\qquad{\rm for}\qquad\texttt{d}=0,1,\ldots,|a|-1\qquad\qquad\big{(}-p-\tfrac{1}{2}\pm{\rm i}s=a\in\mathbb{Z}\,,\ -{\cal S}\leq a<0\big{)}$ (4.43)
@@
${\cal W}_{\rho,\nu}=\bigoplus_{\texttt{d}=0}^{\infty}{\cal H}_{p,s}^{(\texttt{d})}\,,$ (4.45)
@@
$\rho=\mathcal{S}+\frac{1}{2}\,,\qquad\quad\nu=\begin{cases}s-\frac{{\rm i}n}{2}&{\rm for}\ \ \ (-{\rm i}s)>0\\[5.69046pt] s+\frac{{\rm i}n}{2}&{\rm for}\ \ \ (-{\rm i}s)<0\end{cases}\qquad\quad\big{(}\tfrac{n}{2}-{\cal S}\pm{\rm i}s=a\in\mathbb{Z}_{<0}\,,\ -{\cal S}\leq a<0\big{)}\,.$
@@
${\rm ch}_{\rho,\nu}(\texttt{q})=\texttt{q}^{-\frac{1}{12}+\frac{\nu^{2}}{n}+\frac{\rho^{2}}{n+2}}\ \bigg{(}\prod_{m=1}^{\infty}\frac{1}{(1-\texttt{q}^{m})^{2}}\bigg{)}\ \sum_{m=0}^{\infty}(-1)^{m}\,\texttt{q}^{\frac{m^{2}}{2}}\big{(}\,\texttt{q}^{m|\,|\rho|-|\nu|\,|}-\texttt{q}^{(m+1)(|\rho|+|\nu|+1)-\frac{1}{2}}\,\big{)}\,.$ (4.46)
One can check that the dimensions of the level subspaces, obtained by expanding ${\rm ch}_{\rho,\nu}(\texttt{q})$ in a series in q, coincides with the integers $N^{+}_{a}(\texttt{d}|\,{\cal S})$ with $-{\cal S}\leq a<0$ and $\texttt{d}=|a|,\,|a|+1,\,|a|+2,\ldots$.
Finally, we mention that the states $|\psi_{p,s}^{(S^{z})}(\boldsymbol{w})\rangle\in{\cal W}_{\rho,\nu}$ appearing in the scaling limit of the Bethe states, see eq.(4.9), have an important interpretation. They are the simultaneous eigenstates of the family of commuting operators known as the quantum AKNS integrable structure [50, 51]. The function $\mbox{e}^{\frac{{\rm i}}{2}\delta}$ (3.41b) entering into the quantization condition coincides with the eigenvalue of a certain so-called reflection operator [52] computed on $|\psi_{p,s}^{(S^{z})}(\boldsymbol{w})\rangle$, see ref.[45] for details.